suppose an object of mass 2.0 kg is pulled to the right with a force of 20 N. The force makes an angle of 20o with the displacement axis. if the interface is rough with a coefficient of kinetic friction 0.3, what is the net force along the displacement axis?

A) 15 N
B) 16 N
C) 18 N
D) 21 N
Answer #1


Net force F = F1+F2 = 20*cos(20) - (mu_k*N) = 18.8-(0.3*N)

N is the Normal force = mg-F1*sin(20) = (2*9.81)-(20*sin(20)) = 12.77


then F = 18.8-(0.3*12.8) = 14.96 N = 15 N is the final answer